# Vortical waves in a quantum fluid with vector, axial and helical charges. II. Dissipative effects

### v1

@@ 4.6 Generic properties of spectra
$\dfrac{b^{3}}{27a^{2}}\left(\dfrac{1}{\tau_{A}}+\dfrac{1}{\tau_{H}}\right)^{3}-\dfrac{b^{2}\delta}{3ac}\left(\dfrac{1}{\tau_{A}}+\dfrac{1}{\tau_{H}}\right)^{2}+\dfrac{\delta\varsigma}{c}\dfrac{1}{\tau_{A}\tau_{H}}=0\,.$ (130)
From the definitions in Eqs. (3.3) and (127) we can rewrite the previous equation as
$\dfrac{(\sigma_{A}^{\omega})^{3}\left(\frac{1}{\tau_{A}}+\frac{1}{\tau_{H}}\right)^{3}}{\left[(\sigma_{A}^{\omega})^{2}-(\sigma_{H}^{\omega})^{2}\right]}-\dfrac{9(\sigma_{A}^{\omega})^{2}\left(\frac{A^{2}}{\tau_{H}}+\frac{B^{2}}{\tau_{A}}\right)\left(\frac{1}{\tau_{A}}+\frac{1}{\tau_{H}}\right)^{2}}{\left[(A^{2}+B^{2})\sigma_{A}^{\omega}-2AB\sigma_{H}^{\omega}\right]}+\dfrac{27\left[(\sigma_{A}^{\omega})^{2}-(\sigma_{H}^{\omega})^{2}\right]\left(\frac{A^{2}}{\tau_{H}}+\frac{B^{2}}{\tau_{A}}\right)}{\tau_{A}\tau_{H}\left[(A^{2}+B^{2})\sigma_{A}^{\omega}-2AB\sigma_{H}^{\omega}\right]}\\ =0\,,$ (131)
where we have multiplied by $729\left[(\sigma_{A}^{\omega})^{2}-(\sigma_{H}^{\omega})^{2}\right]$ and divided by $\left(\sigma_{A}^{\omega}-\frac{1}{3}T^{2}\Delta H\right)$ to have a more compact expression. Finally, multiplying Eq. (131) by $\tau_{H}^{3}$, it is manifest that it depends on the ratio of lifetimes $\tau_{H}/\tau_{A}\,$.
The solutions to Eq. (131) are shown in Figure 8. The spectra along the blue line feature a point where all three modes coincide, which appear when two branches of propagating modes merge together (see, e.g., Figure 5 in the previous section). We recover the result of the previous section that for a conserved axial charge, i.e., $\tau_{H}/\tau_{A}\to 0$, the merging point exists for $\mu_{V}\simeq 1.4471\,T$. The curve divides the parameter space into two regions: below the curve, there will be two branches of propagating modes, as it happens in the second line of Figure 5; the complementary region contains only one branch of propagating modes, similarly to the last line in the same figure. Furthermore, there is a global maximum at $\mu_{V}/T=0$: $\tau_{A}=2\tau_{H}$. Thus, we can assert that for $\tau_{A}<2\tau_{H}$, there will only be one branch of propagating modes, regardless of the value of the vector chemical potential.
In Figure 9, we show the spectrum of excitations for fixed $\mu_{V}/T=1$ and varying the ratio of axial and helical lifetimes, which can be understood as varying $\tau_{A}$ for fixed $\tau_{H}$. As anticipated from Figure 8, two branches of propagating modes exist for sufficiently small $\tau_{H}/\tau_{A}$. As we increase the ratio, both branches collide, and above a threshold value, there is only one branch of propagating modes. It should also be noted that for a non-conserved axial charge, i.e., finite $\tau_{A}$, there are no propagating modes for small wavenumber.
@@
Refer to caption Refer to caption
Figure 10: Perturbations in the axial chemical potentials at a fixed time $t=5$ fm/c as we vary the axial relaxation time for (left) $c_{h}\tau_{H}/\sigma=100$ (right) $c_{h}\tau_{H}/\sigma=0.1$. The system is initialized with a Gaussian perturbation in the vector chemical potential $\delta\mu_{V}$ on top of a background vector chemical potential $\mu_{V}=650$ MeV at temperature $T=300$ MeV and angular frequency $\Omega=6.6$ MeV.
@@ 5 Dissipative effects
In our considerations so far, we have treated the V/A/H fluid as ideal, ignoring dissipative corrections to the energy-momentum tensor $T^{\mu\nu}$ and the charge currents $J^{\mu}_{\ell}$. In a realistic fluid, dissipation occurs due to the inter-particle interaction, which drives the system towards thermodynamic equilibrium. Contrary to intuition, the perfect fluid limit is achieved when the interactions are sufficiently strong, taking place on time scales much shorter than the characteristic macroscopic time scale, such that equilibration happens almost instantaneously and the deviation of the fluid from thermal equilibrium is always negligible. Since the relaxation time $\tau_{H}$ arises due to helicity-violating pair annihilation (HVPA) two-to-two interactions, the perfect fluid limit implies also that $\tau_{H}\rightarrow 0$. The same limit implies also that the axial relaxation time $\tau_{A}$, describing the dissipation of axial charge, should be taken to 00. In order to provide a more credible assessment of realistic V/A/H fluids, we will consider in this section the effect of dissipation on the spectrum of vortical waves.
@@ 5.1 Relaxation-time approximation
$T^{\mu\nu}\rightarrow T^{\mu\nu}=T^{\mu\nu}_{0}+\pi^{\mu\nu}_{d}-\Pi_{d}\Delta^{\mu\nu},\quad J^{\mu}_{\ell}\rightarrow J^{\mu}_{\ell}=J^{\mu}_{\ell;0}+V^{\mu}_{\ell;d},$ (132)
@@
$\tau_{R}\simeq\frac{T}{P\sigma},$ (134)
@@
The evaluation of transport coefficients in theories with vector interactions is a daunting task, and we do not pursue such calculations in what follows Carrington:2007fp ; Arnold:2006fz . We will consider the qualitative effect of dissipation on the wave spectrum in the frame of kinetic theory under the relaxation-time approximation of Anderson and Witting Anderson:1974nyl . Denoting by $f^{\sigma}_{\mathbf{p},\lambda}$ the number of particles/anti-particles ($\sigma=\pm 1$) with polarization $\lambda$, we postulate a kinetic equation of the form
$p^{\mu}\partial_{\mu}f^{\sigma}_{\mathbf{p},\lambda}=-\frac{E_{\mathbf{p}}\cdot u}{\tau_{R}}\left(f^{\sigma}_{\mathbf{p},\lambda}-f^{{\rm eq};\sigma}_{\mathbf{p},\lambda}\right),$ (135)
@@
$f^{{\rm eq};\sigma}_{\mathbf{p},\lambda}=\left[\exp\left(\frac{p\cdot u-\mu_{\sigma,\lambda}}{T}\right)+1\right]^{-1},$ (136)
is the Fermi-Dirac distribution characterized by the particle four-momentum $p^{\mu}=(p^{0},\mathbf{p})$ with $p^{0}=|\mathbf{p}|$ for massless fermions, $\sigma=\pm 1$ distinguishes between particles and anti-particles, and $\lambda=\pm 1/2$ labels the particle helicity.
Due to its classical nature, Eq. (135) cannot account for the vortical effects that modify the charge currents $J^{\mu}_{\ell}$ in the presence of finite vorticity $\omega^{\mu}$. In order to account for such effects, one must employ a quantum kinetic equation, such as the one derived from the Wigner equation, leading to a modified equilibrium distribution Yang:2020mtz . Similarly, Eq. (135) is unable to account for the non-conservation of the helical and axial charges, thus our current simplified formulation provides access only to the dissipative quantities $V^{\mu}_{\ell;d}$ and $\pi^{\mu\nu}_{d}$.
@@
At equilibrium, when $f^{\sigma}_{\mathbf{p},\lambda}=f^{\sigma;{\rm eq}}_{\mathbf{p},\lambda}$, we have $J^{\mu}_{\ell}=Q_{\ell}u^{\mu}$ and $T^{\mu\nu}=(E+P)u^{\mu}u^{\nu}-Pg^{\mu\nu}$. The conservation equations $\partial_{\mu}J^{\mu}_{\ell}=0$ and $\partial_{\mu}T^{\mu\nu}=0$ imply the Landau matching conditions $u_{\mu}J^{\mu}_{\ell}=Q_{\ell}$ and $u_{\nu}T^{\mu\nu}=Eu^{\mu}$, which we consider to hold in our analysis.
@@ 5.2 Chapman-Enskog expansion
$\delta f^{\sigma}_{\mathbf{p},\lambda}=-\frac{\tau_{R}}{E_{\mathbf{p}}}p^{\mu}\partial_{\mu}f^{\sigma}_{\mathbf{p},\lambda}=-\tau_{R}\left[\partial_{\alpha}\left(\frac{p^{\alpha}}{E_{\mathbf{p}}}f^{\sigma}_{\mathbf{p},\lambda}\right)+\frac{\partial_{\alpha}u_{\beta}}{E_{\mathbf{p}}^{2}}p^{\alpha}p^{\beta}f^{\sigma}_{\mathbf{p},\lambda}\right].$ (138)
@@
$\sum_{\sigma,\lambda}\begin{pmatrix}T^{{\rm eq};\mu\nu\alpha}_{\sigma,\lambda;1}\\ q^{\ell}_{\sigma,\lambda}T^{{\rm eq};\mu\nu\alpha}_{\sigma,\lambda;2}\end{pmatrix}$ $=\frac{1}{3}\begin{pmatrix}E\\ Q_{\ell}\end{pmatrix}(6u^{\mu}u^{\nu}u^{\lambda}-u^{\mu}g^{\nu\lambda}-u^{\nu}g^{\lambda\mu}-u^{\lambda}g^{\mu\nu}),$
$\sum_{\sigma,\lambda}T^{\mu\nu\alpha\beta}_{\sigma,\lambda;2}$ $=\frac{1}{15}E[48u^{\mu}u^{\nu}u^{\lambda}u^{\kappa}-6(u^{\mu}u^{\nu}g^{\lambda\kappa}+u^{\mu}u^{\lambda}g^{\nu\kappa}+u^{\mu}u^{\kappa}g^{\nu\lambda}+u^{\nu}u^{\lambda}g^{\mu\kappa}$
$+u^{\nu}u^{\kappa}g^{\mu\lambda}+u^{\lambda}u^{\kappa}g^{\mu\nu})+g^{\mu\nu}g^{\lambda\kappa}+g^{\mu\lambda}g^{\nu\kappa}+g^{\mu\kappa}g^{\nu\lambda}].$ (141)
@@
$V^{\mu}_{\ell;d}$ $=-\tau_{R}\left[u^{\mu}\partial_{\lambda}(Q_{\ell}u^{\lambda})-\frac{1}{3}\nabla^{\mu}Q_{\ell}+Q_{\ell}Du^{\mu}\right],$
$\pi^{\mu\nu}_{d}$ $=-\tau_{R}\left[(4u^{\mu}u^{\nu}-g^{\mu\nu})DP-u^{\mu}\nabla^{\nu}P-u^{\nu}\nabla^{\mu}P+\frac{24P}{5}(u^{\mu}u^{\nu}\theta+u^{\mu}Du^{\nu}+u^{\nu}Du^{\mu})\right.$
$\left.-\frac{4P}{5}(g^{\mu\nu}\theta-\partial^{\mu}u^{\nu}-\partial^{\nu}u^{\mu})\right],$ (142)
with $Df=u^{\mu}\partial_{\mu}f$ being the comoving derivative and $\nabla^{\mu}f=\Delta^{\mu\nu}\partial_{\nu}f$ the derivative in the fluid’s rest frame.
@@
$DQ_{\ell}+Q_{\ell}\theta$ $=O(\tau_{R}),$ $DE+(E+P)\theta$ $=O(\tau_{R}),$ $(E+P)Du^{\mu}-\nabla^{\mu}P$ $=O(\tau_{R}).$ (143)
@@
where $\sigma^{\mu\nu}=(\frac{1}{2}\Delta^{\mu\lambda}\Delta^{\nu\kappa}+\frac{1}{2}\Delta^{\nu\lambda}\Delta^{\mu\kappa}-\frac{1}{3}\Delta^{\mu\nu}\Delta^{\lambda\kappa})\partial_{\lambda}u_{\kappa}$ is the shear stress tensor, and we identified the shear viscosity as
@@
$V^{\mu}_{\ell;d}=\sum_{\ell^{\prime}}\kappa_{\ell\ell^{\prime}}\nabla^{\mu}\alpha_{\ell^{\prime}},\qquad\kappa_{\ell\ell^{\prime}}=\tau_{R}\left(\frac{1}{3}\frac{\partial Q_{\ell}}{\partial\alpha_{\ell^{\prime}}}-\frac{Q_{\ell}Q_{\ell^{\prime}}T}{E+P}\right).$ (146)

### v2

@@ 4.6 Generic properties of spectra
$\dfrac{b^{3}}{27a^{2}}\left(\dfrac{1}{\tau_{A}}+\dfrac{1}{\tau_{H}}\right)^{3}-\dfrac{b^{2}\delta}{3ac}\left(\dfrac{1}{\tau_{A}}+\dfrac{1}{\tau_{H}}\right)^{2}\\ +\dfrac{\delta\varsigma}{c}\dfrac{1}{\tau_{A}\tau_{H}}=0\,.$ (126)
From the definitions in Eqs. (73) and (123), we can rewrite the previous equation as
$\dfrac{(\sigma_{A}^{\omega})^{3}\left(\frac{1}{\tau_{A}}+\frac{1}{\tau_{H}}\right)^{3}}{\left[(\sigma_{A}^{\omega})^{2}-(\sigma_{H}^{\omega})^{2}\right]}\\ +\dfrac{27\left[(\sigma_{A}^{\omega})^{2}-(\sigma_{H}^{\omega})^{2}\right]\left(\frac{A^{2}}{\tau_{H}}+\frac{B^{2}}{\tau_{A}}\right)}{\tau_{A}\tau_{H}\left[(A^{2}+B^{2})\sigma_{A}^{\omega}-2AB\sigma_{H}^{\omega}\right]}\\ -\dfrac{9(\sigma_{A}^{\omega})^{2}\left(\frac{A^{2}}{\tau_{H}}+\frac{B^{2}}{\tau_{A}}\right)\left(\frac{1}{\tau_{A}}+\frac{1}{\tau_{H}}\right)^{2}}{\left[(A^{2}+B^{2})\sigma_{A}^{\omega}-2AB\sigma_{H}^{\omega}\right]}=0\,,$ (127)
where we multiplied by $729\left[(\sigma_{A}^{\omega})^{2}-(\sigma_{H}^{\omega})^{2}\right]$ and divided by $\left(\sigma_{A}^{\omega}-\frac{1}{3}T^{2}\Delta H\right)$ to have a more compact expression. Finally, multiplying Eq. (127) by $\tau_{H}^{3}$, it is manifest that it depends on the ratio of lifetimes $\tau_{H}/\tau_{A}\,$.
The solutions of Eq. (127) are shown in Figure 8. The spectra along the blue line feature a point where all three modes coincide, which appear when two branches of propagating modes merge together (see, e.g., Figure 5 in the previous section). We recover the result of the previous section that for a conserved axial charge, i.e., $\tau_{H}/\tau_{A}\to 0$, the merging point exists for $\mu_{V}\simeq 1.4471\,T$. The curve divides the parameter space into two regions: below the curve, there will be two branches of propagating modes, as it happens in the second line of Figure 9; the complementary region contains only one branch of propagating modes, similarly to the last line in Fig. 9. Furthermore, the curve in Fig. 8 has a global maximum at $\mu_{V}/T=0$: $\tau_{A}=2\tau_{H}$. Thus, we can assert that for $\tau_{A}<2\tau_{H}$, there will only be one branch of propagating modes, regardless of the value of the vector chemical potential.
In Fig. 9, we show the spectrum of excitations for fixed $\mu_{V}/T=1$ and varying the ratio of axial and helical lifetimes, which can be understood as varying $\tau_{A}$ for fixed $\tau_{H}$. As anticipated from Fig. 8, two branches of propagating modes exist for sufficiently small $\tau_{H}/\tau_{A}$. As we increase the ratio, these branches collide, and above a threshold value, there is only one branch of propagating modes. It should also be noted that for a non-conserved axial charge, i.e., finite $\tau_{A}$, there are no propagating modes for small wavenumber.
@@
Figure 10: Perturbations in the axial chemical potentials at a fixed time, $c_{h}t/\sigma=5$, as we vary the axial relaxation time for (top) $c_{h}\tau_{H}/\sigma=100$ (bottom) $c_{h}\tau_{H}/\sigma=0.1$. The system is initialized with a Gaussian perturbation in the vector chemical potential $\delta\mu_{V}$ on top of a background dimensionless vector chemical potential $\alpha_{V}=13/6$.
@@ 5 Dissipative effects
In our considerations so far, we have treated the V/A/H fluid as ideal, ignoring dissipative corrections to the energy-momentum tensor $T^{\mu\nu}$ and the charge currents $J^{\mu}_{\ell}$. In a realistic fluid, dissipation occurs due to the inter-particle interaction, which drives the system towards thermodynamic equilibrium. Contrary to intuition, the perfect fluid limit is achieved when the interactions are sufficiently strong, taking place on time scales much shorter than the characteristic macroscopic time scale, such that equilibration happens almost instantaneously and the deviation of the fluid from thermal equilibrium is always negligible. Since the relaxation time $\tau_{H}$ arises due to helicity-violating pair annihilation (HVPA) two-to-two interactions, the perfect fluid limit implies also that $\tau_{H}\rightarrow 0$. Similarly, the same limit implies also that the axial relaxation time $\tau_{A}$, describing the dissipation of axial charge, should be taken to 00. In order to provide a more credible assessment of realistic V/A/H fluids, we will consider in this section the effect of dissipation on the spectrum of vortical waves.
@@ 5.1 Relaxation-time approximation
$T^{\mu\nu}\rightarrow T^{\mu\nu}=T^{\mu\nu}_{0}+\pi^{\mu\nu}_{d}-\Pi_{d}\Delta^{\mu\nu},$
$J^{\mu}_{\ell}\rightarrow J^{\mu}_{\ell}=J^{\mu}_{\ell;0}+V^{\mu}_{\ell;d},$ (128)
@@
$\tau_{R}\propto\frac{T}{P\sigma},$ (130)
@@
The evaluation of transport coefficients in theories with vector interactions is a daunting task Carrington:2007fp ; Arnold:2006fz , and we do not pursue such calculations in what follows. We will consider the qualitative effect of dissipation on the wave spectrum in the frame of kinetic theory under the relaxation-time approximation of Anderson and Witting Anderson:1974nyl . Denoting by $f^{\sigma}_{\mathbf{p},\lambda}$ the number of particles/anti-particles ($\sigma=\pm 1$) with polarization $\lambda=\pm 1/2$, we postulate a kinetic equation of the form
$p^{\mu}\partial_{\mu}f^{\sigma}_{\mathbf{p},\lambda}=-\frac{E_{\mathbf{p}}\cdot u}{\tau_{R}}\left(f^{\sigma}_{\mathbf{p},\lambda}-f^{\sigma;{\rm eq}}_{\mathbf{p},\lambda}\right),$ (131)
@@
$f^{\sigma{\rm eq}}_{\mathbf{p},\lambda}=\left[\exp\left(\frac{p\cdot u-\mu_{\sigma,\lambda}}{T}\right)+1\right]^{-1},$ (132)
is the Fermi-Dirac distribution characterized by the particle four-momentum $p^{\mu}=(p^{0},\mathbf{p})$ with $p^{0}=\lvert\mathbf{p}\rvert$ for massless fermions, $\sigma=\pm 1$ distinguishes between particles and anti-particles, and $\lambda=\pm 1/2$ labels the particle helicity.
Due to its classical nature, Eq. (131) cannot account for the vortical effects that modify the charge currents $J^{\mu}_{\ell}$ in the presence of finite vorticity $\omega^{\mu}$. In order to account for such effects, one must employ a quantum kinetic equation, such as the one derived from the Wigner equation, leading to a modified equilibrium distribution Yang:2020mtz . An alternative approach, followed in Ref. Gorbar:2017toh in the absence of helicity, is to make use of Chiral Kinetic Theory to describe the vorticity contributions to the currents. Similarly, Eq. (131) is unable to account for the non-conservation of the helical and axial charges, thus our current simplified formulation provides access only to the dissipative quantities $V^{\mu}_{\ell;d}$ and $\pi^{\mu\nu}_{d}$.
@@
[…]
At equilibrium, when $f^{\sigma}_{\mathbf{p},\lambda}=f^{\sigma;{\rm eq}}_{\mathbf{p},\lambda}$, we have $J^{\mu}_{\ell}=Q_{\ell}u^{\mu}$ and $T^{\mu\nu}=(E+P)u^{\mu}u^{\nu}-Pg^{\mu\nu}$. The conservation equations $\partial_{\mu}J^{\mu}_{\ell}=0$ and $\partial_{\mu}T^{\mu\nu}=0$ imply the Landau matching conditions $u_{\mu}J^{\mu}_{\ell}=Q_{\ell}$ and $u_{\nu}T^{\mu\nu}=Eu^{\mu}$, valid also when the fluid is out of equilibrium, which we consider to hold in our analysis.
@@ 5.2 Chapman-Enskog expansion
$\delta f^{\sigma}_{\mathbf{p},\lambda}=-\frac{\tau_{R}}{E_{\mathbf{p}}}p^{\mu}\partial_{\mu}f^{\sigma}_{\mathbf{p},\lambda}$
$=-\tau_{R}\left[\partial_{\alpha}\left(\frac{p^{\alpha}}{E_{\mathbf{p}}}f^{\sigma}_{\mathbf{p},\lambda}\right)+\frac{\partial_{\alpha}u_{\beta}}{E_{\mathbf{p}}^{2}}p^{\alpha}p^{\beta}f^{\sigma}_{\mathbf{p},\lambda}\right].$ (134)
@@
$\sum_{\sigma,\lambda}\begin{pmatrix}T^{{\rm eq};\mu\nu\alpha}_{\sigma,\lambda;1}\\ q^{\ell}_{\sigma,\lambda}T^{{\rm eq};\mu\nu\alpha}_{\sigma,\lambda;2}\end{pmatrix}=\frac{1}{3}\begin{pmatrix}E\\ Q_{\ell}\end{pmatrix}(6u^{\mu}u^{\nu}u^{\alpha}$
$\hskip 86.72267pt-u^{\mu}g^{\nu\alpha}-u^{\nu}g^{\alpha\mu}-u^{\alpha}g^{\mu\nu}),$
$\sum_{\sigma,\lambda}T^{\mu\nu\alpha\beta}_{\sigma,\lambda;2}=\frac{1}{15}E[48u^{\mu}u^{\nu}u^{\alpha}u^{\beta}$
$-6(u^{\mu}u^{\nu}g^{\alpha\beta}+u^{\mu}u^{\alpha}g^{\nu\beta}+u^{\mu}u^{\beta}g^{\nu\alpha}+u^{\nu}u^{\alpha}g^{\mu\beta}$
$+u^{\nu}u^{\beta}g^{\mu\alpha}+u^{\alpha}u^{\beta}g^{\mu\nu})+g^{\mu\nu}g^{\alpha\beta}+g^{\mu\alpha}g^{\nu\beta}$
$\hskip 260.17464pt+g^{\mu\beta}g^{\nu\alpha}].$ (137)
@@
$V^{\mu}_{\ell;d}$ $=-\tau_{R}\left[u^{\mu}\partial_{\alpha}(Q_{\ell}u^{\alpha})-\frac{1}{3}\nabla^{\mu}Q_{\ell}+Q_{\ell}Du^{\mu}\right],$
$\pi^{\mu\nu}_{d}$ $=-\tau_{R}\left[(3u^{\mu}u^{\nu}-\Delta^{\mu\nu})DP-u^{\mu}\nabla^{\nu}P-u^{\nu}\nabla^{\mu}P\right.$
$\left.+4P\left(\frac{6}{5}u^{\mu}u^{\nu}\theta+u^{\mu}Du^{\nu}+u^{\nu}Du^{\mu}\right)\right.$
$\left.-\frac{4P}{5}(g^{\mu\nu}\theta+\nabla^{\mu}u^{\nu}+\nabla^{\nu}u^{\mu})\right],$ (138)
with $Df=u^{\mu}\partial_{\mu}f$ and $\nabla^{\mu}f=\Delta^{\mu\nu}\partial_{\nu}f$.
@@
$DQ_{\ell}+Q_{\ell}\theta$ $=O(\tau_{R}),$
$DE+(E+P)\theta$ $=O(\tau_{R}),$
$(E+P)Du^{\mu}-\nabla^{\mu}P$ $=O(\tau_{R}).$ (139)
@@
where $\sigma^{\mu\nu}=(\frac{1}{2}\Delta^{\mu\lambda}\Delta^{\nu\kappa}+\frac{1}{2}\Delta^{\nu\lambda}\Delta^{\mu\kappa}-\frac{1}{3}\Delta^{\mu\nu}\Delta^{\lambda\kappa})\partial_{\lambda}u_{\kappa}=\frac{1}{2}(\nabla^{\mu}u^{\nu}+\nabla^{\nu}u^{\mu})-\frac{1}{3}\Delta^{\mu\nu}\theta$ is the shear stress tensor, and we identified the shear viscosity as
@@
$V^{\mu}_{\ell;d}=\sum_{\ell^{\prime}}\kappa_{\ell\ell^{\prime}}\nabla^{\mu}\alpha_{\ell^{\prime}},$
$\kappa_{\ell\ell^{\prime}}=\tau_{R}\left(\frac{1}{3}\frac{\partial Q_{\ell}}{\partial\alpha_{\ell^{\prime}}}-\frac{Q_{\ell}Q_{\ell^{\prime}}T}{E+P}\right).$ (142)